Calculate the quotient [HCO2-]/[HCO2H] at pH 3.744 if the ionic strength is 0.1M by using the effective equilibrium constant listed for ionic strength = 0.1

Please show all work and explain. Thanks!
Answer #1

Balanced equilibirium:
HCOOH <----> HCOO- + H+
Use Henderson Hasselbalch equation:
pH = pKa + log([HCOO-]/[HCOOH])
Given, pH = 3.744
Ka = 0.1
so, pKa = -log0.1 = 1
Rearrange the equation:
log([HCOO-]/[HCOOH])    = pH-pKa
= 3.744-1
= 2.744
([HCOO-]/[HCOOH])       = 10^2.744
= 554.6
Answer #2

The equilibrium constant for the reaction